# Are Ensembles Getting Better All the Time?

### v2

@@ Abstract
Ensemble methods combine the predictions of several base models. We study whether or not including more models always improves their average performance. This question depends on the kind of ensemble considered, as well as the predictive metric chosen. We focus on situations where all members of the ensemble are a priori expected to perform as well, which is the case of several popular methods such as random forests or deep ensembles. In this setting, we show that ensembles are getting better all the time if, and only if, the considered loss function is convex. More precisely, in that case, the average loss of the ensemble is a decreasing function of the number of models. When the loss function is nonconvex, we show a series of results that can be summarised as: ensembles of good models keep getting better, and ensembles of bad models keep getting worse. To this end, we prove a new result on the monotonicity of tail probabilities that may be of independent interest. We illustrate our results on a medical prediction problem (diagnosing melanomas using neural nets) and a “wisdom of crowds” experiment (guessing the ratings of upcoming movies).
@@ 1 Introduction
The empirical successes of ensembles seem to indicate that the more models are being aggregated, the better the ensemble is. The theoretical question we address in this paper is what we call the monotonicity question: Is it always true that an ensemble of $K+1$ models performs better than an ensemble of $K$ models? A positive answer to this question has practical consequences: given an ensemble with $K$ base models, performances will improve when using one (or ten) more.
Figure 1: Evolution of the accuracy of a dropout ensemble on the DermaMNIST dataset as the number of models grows (mean and standard deviation over $500$ repetitions). (Left) The accuracy averaged over the whole test set has no clear monotonic pattern. (Middle) The accuracy averaged over images for which the asymptotic prediction $\overline{y}_{\infty}$ is correct is getting better. (Right) The accuracy averaged over images for which the asymptotic prediction $\overline{y}_{\infty}$ in incorrect is getting worse. The behaviours of these middle and right panels are explained by our theory. Notice that the three $y$-axes have different scales.
@@ 1.1 A motivating example: neural nets for medical diagnosis
In this section, we consider the following practical problem: we want to predict whether a dermatoscopic image corresponds to a benign keratosis-like lesion (good) or a melanoma (bad). We use the DermaMNIST (Yang et al., 2023) data set, based on the HAM10000 collection (Tschandl et al., 2018), and retain only the classes “benign keratosis” and “melanoma” (more details on the data set are available in Section 5). Using a training set and a validation set, we train a LeNet-like convolutional neural network (LeCun et al., 1998) with dropout to discriminate these two classes. We then use a Monte Carlo dropout ensemble to classify the test set (Gal and Ghahramani, 2016; Hron et al., 2018), and look at the evolution of the accuracy (Figure 1) and the cross-entropy loss (Figure 2) as the ensemble grows. There is a crisp difference between the behaviours of the two losses. The overall accuracy has no clear monotonic pattern, but if we split the test set into two subsets depending on whether the asymptotic prediction is right or wrong, we notice that the accuracy is decreasing for right predictions and increasing for wrong ones! On the other hand, the cross entropy appears to be clearly decreasing, regardless of the quality of the prediction. Our goal is to provide clear theoretical explanations for these phenomena.
Figure 2: Evolution of the cross-entropy (or negative log-likelihood) of a dropout ensemble on the dermatology data set as the number of models grows (mean and standard deviation over 500 repetitions). (Left) Cross-entropy averaged over the whole test set. (Middle) Cross-entropy averaged over images for which the asymptotic prediction $\overline{y}_{\infty}$ is correct. (Right) Cross-entropy averaged over images for which the asymptotic prediction $\overline{y}_{\infty}$ in incorrect. As predicted by our theory, all curves are decreasing. Notice that the three $y$-axes have different scales.
@@ 1.2 Summary and organisation of the paper
In this work, we focus on the simplest instance of ensembles, that involve unweighted averaging of predictions that are “not too dependent.” While simple, this setting remains the usual formalization of most “wisdom of crowds” studies, and the basis of state-of-the-art machine learning methods like random forests and deep ensembles. In this context, introduced in Section 2, for an ensemble to be monotonically improving is deeply connected to the local geometry of the chosen loss function, more specifically to its convexity and its smoothness. More precisely, we show that:
@@
  • For nonconvex loss functions and independent and identically distributed (i.i.d.) predictions (Section 4), we show that, asymptotically, good ensembles keep getting better, and bad ensembles keep getting worse in the following sense:
@@ 1.3 Related works: monotonicity and the wisdom of crowds
The main heuristic justification of ensembles is that groups are collectively wiser than individuals, because aggregated decisions have a smaller variance than individual ones. This general idea is referred to in various, sometimes colorful, ways: some, like Dekel and Shamir (2009), praise the “vox populi,” following Galton’s (1907) seminal paper; the phrase “wisdom of crowds,” used for instance by Prelec et al. (2017) or Simoiu et al. (2019), was immensely popularised by Surowiecki’s (2005) bestselling book; political scientists like Althaus (2003) often debate the virtues of the “miracle of aggregation” as a theoretical argument in favor of democracy; many scientists made the popular phrase “two heads are better than one” their own (e.g., Koriat, 2012; Chu-Carroll et al., 2003).
@@
Condorcet (1785) proposed a first mathematical formalisation of this heuristic insight. His goal was to prove that, within a trial, the larger the number of jurors, the better. In modern machine learning jargon, he showed that ensembles containing an odd number of Bernoulli-distributed predictors are monotonically improving, with respect to the classification error (see Shteingart et al. 2020, for a modern proof). Surprisingly, he also noticed that monotonicity is broken if the ensemble is allowed to contain an even number of predictors.
@@
While an important amount of work has been devoted to the theoretical study of ensembles (see, e.g., Germain et al., 2015; Scornet et al., 2015; Lopes, 2019; Ortega et al., 2022; Larsen, 2023), the monotonicity question has been quite overlooked, especially theoretically. Lam and Suen (1997) and Dougherty and Edward (2009) revisited Condorcet’s theorem, and studied the issues that make the presence of an even number of models problematic. Several generalisations of this theorem were also derived, in particular with weaker dependence assumptions (see, e.g., Ladha, 1993). Probst and Boulesteix (2018), in the context of random forests, performed a very thorough theoretical and empirical study of the monotonicity question. Their empirical conclusions were that, for some performance metrics (the cross-entropy or the Brier score for classification and the mean squared error for regression), ensembles are getting better all the time, but that is not necessarily the case for the classification error. Moreover, they provided theoretical support for some of these phenomena. Their paper was a key inspiration for this work, and our main contribution is to provide a unified view of their insights by essentially showing that ensembles are getting better all the time if, only if, the considered loss function is convex. Indeed, the cross-entropy, the Brier score, and the mean squared error are all convex losses (and always lead to monotonic improvements), whereas the classification error is not (and may lead to non-monotonicity).
@@
It is worth mentioning some related work on monotonicity in statistics and machine learning. A first intriguing phenomenon is the fact that the coverage probability of a confidence interval for the true parameter of a Bernoulli random variable is non-monotonic (see, e.g., Brown et al. 2001, Figure 1). In plain words, and perhaps counter-intuitively, increasing the number of observations does not guarantee that the probability for the confidence interval to contain the true parameter increases! Another related question is the one of the monotonicity of learning: will a machine learning algorithm perform better when adding one new sample to the training set? This problem is mentioned for instance by Rasmussen and Williams (2006, Exercise 2.4), and was recently cast as an open problem by Viering et al. (2019) at the 32nd Conference on Learning Theory (COLT). Progress on that front was recently reviewed by Viering and Loog (2023).
@@ 2.1 A general predictive framework
We consider a predictive task where several methods output predictions $\hat{y}$, living in a convex set $C$. The quality of these predictions is evaluated using a loss function $L:C\rightarrow\mathbb{R}$. This broad setting contains several important cases, and we present below a few important examples. In most of these examples, $\hat{y}$ will be of the form $f(x)$, where $f$ is a predictor and $x$ a given data point. This means that $L(\hat{y})$ is the loss of an individual prediction, rather than the loss averaged over a dataset. Note that we require the prediction set to be convex because we want that averaged predictions remain inside $C$ (this is barely a restriction, since when the prediction set is nonconvex we can simply replace it by its convex hull).
@@
where the true label is represented by a one-hot encoding $y=(y_{1},\ldots,y_{n_{\text{Cl}}})\in\{0,1\}^{n_{\text{Cl}}}$. Another common choice of loss function in that case is the classification error, also known as the $0-1$ loss,
$L(\hat{y})=\ell(\hat{y},y)=\left\{\begin{matrix}1\;\text{ if }\;\text{argmax}(\hat{y})\neq\text{argmax}(y)\\ 0\;\text{ if }\;\text{argmax}(\hat{y})=\text{argmax}(y)\,.\end{matrix}\right.$ (2)
If the response set is continuous (e.g., if $\mathcal{Y}=\mathbb{R}$ as in a standard regression problem), the prediction set can be simply $\mathbb{R}$ if we are simply interested in predicting a point estimate of the response, and $\ell$ can be the squared error: $L(\hat{y})=\ell(\hat{y},y)=(\hat{y}-y)^{2}$. In we are interested in the full conditional distribution of the response, $C$ can be defined as the set of probability densities over $\mathbb{R}$. The loss $\ell$ can then be any scoring rule like the negative log-likelihood $-\log\hat{y}(y|x)$ or the continuous ranked probability score, see, e.g., Gneiting and Raftery 2007. Other supervised examples with more exotic prediction spaces and loss functions include include image segmentation or ordinal regression.
Density estimation is generally unsupervised, and not directly a predictive problem. It is nevertheless a particular case of our framework. Here, $\hat{y}$ will be a probability density over the data space $\mathcal{X}$, for instance modelled by a mixture model or a deep generative model, which means that $C$ will be the set of all densities over $\mathcal{X}$. For a given $x\in\mathcal{X}$, a popular loss function is then the negative log-likelihood $L(\hat{y})=-\log\hat{y}(x)$. Similarly to the probabilistic regression example outlined above, any scoring rule can also be chosen (e.g., score matching and its variants, commonly used for density estimation). More exotic losses include using statistical divergences between the distribution $\hat{y}$ and the empirical distribution of a test or validation set. While ensembling is less common for density estimation than forecasting and supervised learning, it has been successfully used in several contexts. For instance, Riesselman et al. (2018) used ensembles of deep generative models to model biological sequences and Russell et al. (2015) and Casa et al. (2021) have used ensembles of Gaussian mixtures for density estimation and clustering.
@@
In many of these examples, the loss function depends on other quantities than the predictions (the features, the true label…), yet we just denote it by $L(\hat{y})$. Our analysis aims at studying the impact of the predictions on the loss, therefore we keep all these other quantities fixed, and vary only the predictions. This is why we only model our loss function $L$ as a function of the predictions $\hat{y}$. In particular, this means that our results typically concern the loss evaluated for a single given data point and a single ground truth label. Of course, we are also often interested in evaluating the loss not only on individual data points, but on batches or entire data sets. Our results on individual data points can then be extended by averaging them, as we discuss further in Section 2.2.3.
@@ 2.2.1 Where do our predictions $\hat{y}_{1},\ldots,\hat{y}_{K}$ come from?
In all the examples presented in the previous paragraph, the $\hat{y}_{1},\ldots,\hat{y}_{K}$ are i.i.d. random variables. Assumptions of this sort will be required for our proofs, although we manage to prove our results under weak dependence structures. We refer to Section 3.3 for a simple example of what can go wrong when there is too much dependency between the individual models.
We want to emphasise that our framework also encompasses situations where we are given a few predictions that are fully deterministic, for instance, the forecasts of several different physical models. One way to see them as random predictions and to make them amenable to our theory is to assume that the ordering of the $\hat{y}_{1},\ldots,\hat{y}_{K}$ does not matter, or equivalently that we have no knowledge of how these models fare against each other. This can be formalised by assuming that we do not observe $\hat{y}_{1},\ldots,\hat{y}_{K}$ but $\hat{y}_{\sigma(1)},\ldots,\hat{y}_{\sigma(K)}$, where $\sigma$ is a uniformly sampled permutation of $\{1,\ldots,K\}$. Because $\sigma$ is random, $\hat{y}_{\sigma(1)},\ldots,\hat{y}_{\sigma(K)}$ is also random. While not i.i.d. in general, this randomly reordered ensemble satisfies another weaker condition called exchangeability. A sequence of random variables is exchangeable when any of its reorderings has the same distribution, we will give a formal definition in Eq. (38).

### v3

@@ Abstract
Ensemble methods combine the predictions of several base models. We study whether or not including more models always improves their average performance. This question depends on the kind of ensemble considered, as well as the predictive metric chosen. We focus on situations where all members of the ensemble are a priori expected to perform equally well, which is the case of several popular methods such as random forests or deep ensembles. In this setting, we show that ensembles are getting better all the time if, and only if, the considered loss function is convex. More precisely, in that case, the loss of the ensemble is a decreasing function of the number of models. When the loss function is nonconvex, we show a series of results that can be summarised as: ensembles of good models keep getting better, and ensembles of bad models keep getting worse. To this end, we prove a new result on the monotonicity of tail probabilities that may be of independent interest. We illustrate our results on a medical problem (diagnosing melanomas using neural nets) and a “wisdom of crowds” experiment (guessing the ratings of upcoming movies).
@@ 1 Introduction
The empirical successes of ensembles seem to indicate that the more models are being aggregated, the better the ensemble is. This, of course, can only be true on average. Adding a consistently bad model to the family will decrease the overall accuracy—for instance adding a model that always predicts the wrong class. The theoretical question we address in this paper is the following: Is it always true that an ensemble of $K+1$ models performs better on average than an ensemble of $K$ models? We refer to this question as the monotonicity question. A positive answer would have practical consequences, encouraging us to always take more base models in our ensemble. As we will see, the picture is more nuanced and needs precise analysis to emerge.
Figure 1: Evolution of the accuracy of a dropout ensemble on the DermaMNIST dataset as the number of models grows (mean and standard deviation over $500$ repetitions). (Left) The accuracy averaged over the whole test set has no clear monotonic pattern. (Middle) Accuracy averaged over images for which the asymptotic prediction (i.e. the prediction of an ensemble of infinite size) $\overline{y}_{\infty}$ is correct. The accuracy is getting better when the ensemble size grows. (Right) Accuracy averaged over images for which $\overline{y}_{\infty}$ is incorrect. The accuracy is going down. The behaviours of these middle and right panels are related to the non-convexity of the classfication error, and are explained by our theory. Notice that the three $y$-axes have different scales.
@@ 1.1 A motivating example: neural nets for medical diagnosis
In this section, we consider the following practical problem: we want to predict whether a dermatoscopic image corresponds to a benign keratosis-like lesion or a melanoma. We use the DermaMNIST (Yang et al., 2023) data set, based on the HAM10000 collection (Tschandl et al., 2018), and retain only the classes “benign keratosis” and “melanoma” (more details on the data set are available in Section 5). Using a training set and a validation set, we train a LeNet-like convolutional neural network (LeCun et al., 1998) with dropout to discriminate these two classes. We then use a Monte Carlo dropout ensemble to classify the test set (Gal and Ghahramani, 2016; Hron et al., 2018), and look at the evolution of the accuracy (Figure 1) and the cross-entropy loss (Figure 2) as the ensemble grows. There is a crisp difference between the behaviours of the two losses. The overall accuracy has no clear monotonic pattern, but if we split the test set into two subsets depending on whether the asymptotic prediction is right or wrong, we notice that the accuracy is increasing for right predictions and decreasing for wrong ones! On the other hand, the cross-entropy appears to be clearly decreasing, regardless of the quality of the prediction. Our goal is to provide clear theoretical explanations for these phenomena.
Figure 2: Evolution of the cross-entropy (or negative log-likelihood) of a dropout ensemble on the dermatology data set as the number of models grows (mean and standard deviation over 500 repetitions). (Left) Cross-entropy averaged over the whole test set. (Middle) Cross-entropy averaged over images for which the asymptotic prediction $\overline{y}_{\infty}$ is correct. (Right) Cross-entropy averaged over images for which the asymptotic prediction $\overline{y}_{\infty}$ in incorrect. As predicted by our theory, all curves are decreasing because the cross-entropy loss is convex. Notice that the three $y$-axes have different scales.
@@ 1.2 Summary and organisation of the paper
In this work, we focus on the simplest instance of ensembles, that involve unweighted averaging of predictions that are “not too dependent.” While simple, this setting remains the usual formalization of most “wisdom of crowds” studies, and the basis of state-of-the-art machine learning methods like random forests and deep ensembles. In this context, formalised in Section 2, for an ensemble to be monotonically improving is deeply connected to the local geometry of the chosen loss function, more specifically to its convexity and its smoothness. More precisely, we show that:
@@
  • For nonconvex loss functions and independent and identically distributed (i.i.d.) predictions (Section 4), we show that, when the number of base models is large enough, good ensembles keep getting better, and bad ensembles keep getting worse in the following sense:
@@ 1.3 Related works: monotonicity and the wisdom of crowds
The main heuristic justification of ensembles is that groups are collectively wiser than individuals, because aggregated decisions have a smaller variance than individual ones. This general idea is referred to in various, sometimes colorful, ways: some, like Dekel and Shamir (2009), praise the “vox populi,” following Galton’s (1907) seminal paper; the phrase “wisdom of crowds,” used for instance by Prelec et al. (2017) or Simoiu et al. (2019), was immensely popularised by Surowiecki’s (2004) bestselling book; political scientists like Althaus (2003) often debate the virtues of the “miracle of aggregation” as a theoretical argument in favor of democracy; many scientists made the popular phrase “two heads are better than one” their own (e.g., Koriat, 2012; Chu-Carroll et al., 2003). A recent review of the philosophical ramifications of collective intelligence was provided by Elkin and Pettigrew (2025).
@@
Condorcet (1785) proposed a first mathematical formalisation of this heuristic insight. His goal was to prove that, within a trial, the larger the number of jurors, the better. In modern machine learning jargon, he showed that ensembles containing an odd number of Bernoulli-distributed base models are monotonically improving, with respect to the classification error (see Shteingart et al. 2020, for a modern proof). Surprisingly, he also noticed that monotonicity is broken if the ensemble is allowed to contain an even number of models.
An important (and quite diverse) amount of work has been devoted to the theoretical study of ensembles; while giving a detailed overview is out of our scope, it is worth providing a few entry points to this literature. Some authors analysed the properties of specific kinds of ensembles, such as bagging (Larsen, 2023), random forests (Scornet and Hooker, 2025), random projections (Cannings and Samworth, 2017), dropout (Hron et al., 2018) or deep ensembles (Wild et al., 2023). Others studied the relationship between the benefits of ensembling and the diversity of the ensemble (Ortega et al., 2022; Theisen et al., 2023). More related to our work are papers that study the behaviour of ensembles when $K$ is large. In particular, Cannings and Samworth (2017) and Lopes (2020) derived asymptotic expansions of the test loss when $K\rightarrow\infty$. Unfortunately, such expansions do not capture monotonicity, even for large enough $K$ (this will be discussed further in Section 4.2).
@@
However, in spite of this large body of work, the monotonicity question has been quite overlooked, especially theoretically. Lam and Suen (1997), Berend and Sapir (2005), and Dougherty and Edward (2009) did revisit Condorcet’s theorem, and studied the issues that make the presence of an even number of models problematic. Several generalisations of this theorem were also derived, in particular with weaker dependence assumptions (see, e.g., Ladha, 1993). Closer to our investigations Probst and Boulesteix (2018), in the context of random forests, performed a very thorough theoretical and empirical study of the monotonicity question. Their empirical conclusions were that, for some performance metrics (the cross-entropy or the Brier score for classification and the mean squared error for regression), ensembles are getting better all the time, but that is not necessarily the case for the classification error. Moreover, they provided theoretical support for some of these phenomena. Their paper was a key inspiration for this work, and our main contribution is to provide a unified view of their insights by essentially showing that ensembles are getting better all the time if, only if, the considered loss function is convex. Indeed, the cross-entropy, the Brier score, and the mean squared error are all convex losses (and always lead to monotonic improvements), whereas the classification error is not (and may lead to non-monotonicity).
@@
It is worth mentioning some related work on monotonicity in statistics and machine learning. A first intriguing phenomenon is the fact that the coverage probability of a confidence interval for the true parameter of a Bernoulli random variable is non-monotonic (see, e.g., Brown et al. 2001, Figure 1). In plain words, and perhaps counter-intuitively, increasing the number of observations does not guarantee that the probability for the confidence interval to contain the true parameter increases! Another related question is the one of the monotonicity of learning: will a machine learning algorithm perform better when adding one new sample to the training set? Devroye et al. (2013) call such class of models smart rules predictors and notices that it is quite challenging to know for sure which models are “smart.” Indeed, most consistency results only provide an asymptotic statement in the sample size, with no monotonicity guarantee. More recently, this problem was cast as an open problem by Viering et al. (2019) at the 32nd Conference on Learning Theory (COLT). Progress on that front was reviewed by Viering and Loog (2023). Finally, the question of monotonicity also arises when looking at the variance of a model, as illustrated, for instance by Rasmussen and Williams (2006, Exercise 2.4), for Gaussian processes.
@@ 2.1 A general predictive framework
We consider a predictive task where several methods output predictions $\hat{y}$, living in a convex set $C$. The quality of these predictions is evaluated using a loss function $L:C\rightarrow\mathbb{R}$. This broad setting contains several important cases, and we present below a few important examples. In most of these examples, $\hat{y}$ will be of the form $f(x)$, where $f$ is a base model and $x$ a given fixed data point. This means that L(y^)L($\hat${y}) is the loss of an individual prediction, rather than the loss averaged over a dataset. Note that we require the prediction set to be convex because we want that averaged predictions remain inside $C$ (this is barely a restriction, since when the prediction set is nonconvex we can simply replace it by its convex hull).
@@
where the true label is represented by a one-hot encoding $y=(y_{1},\ldots,y_{n_{\text{Cl}}})\in\{0,1\}^{n_{\text{Cl}}}$. Another common choice of loss function in that case is the classification error, also known as the 010-1 loss. Denoting $\text{argmax}(u)$ the index of the coordinate of $u$ with the largest value, the classification error can be written
[…]
If the response set is continuous (e.g., if $\mathcal{Y}=\mathbb{R}$ as in a standard regression problem), the prediction set can be $\mathbb{R}$ if we are simply interested in predicting a point estimate of the response, then $\ell$ can be the squared error: $L(\hat{y})=\ell(\hat{y},y)=(\hat{y}-y)^{2}$. If we are interested in the full conditional distribution of the response, $C$ can be defined as the set of probability densities over $\mathbb{R}$. The loss $\ell$ can then be any scoring rule like the negative log-likelihood $\log\hat{y}(y|x)$ or the continuous ranked probability score, see, e.g., Gneiting and Raftery (2007). Other supervised examples with more exotic prediction spaces and loss functions include image segmentation or ordinal regression.
Density estimation is generally unsupervised, and not directly a predictive problem. It is nevertheless a particular case of our framework. Here, $\hat{y}$ will be a probability density over the data space $\mathcal{X}$, for instance modelled by a mixture model or a deep generative model, which means that $C$ will be the set of all densities over $\mathcal{X}$. For a given $x\in\mathcal{X}$, a popular loss function is then the negative log-likelihood $L(\hat{y})=-\log\hat{y}(x)$. Similarly to the probabilistic regression example outlined above, any scoring rule can also be chosen (e.g., score matching and its variants). More exotic losses include using statistical divergences between the distribution $\hat{y}$ and the empirical distribution of a test or validation set. While ensembling is less common for density estimation than forecasting and supervised learning, it has been successfully used in several contexts. For instance, Riesselman et al. (2018) used ensembles of deep generative models to model biological sequences and Russell et al. (2015) and Casa et al. (2021) have used ensembles of Gaussian mixtures for density estimation and clustering.
@@
In many of these examples, the loss function depends on other quantities than the predictions (the features, the true label…), yet we just denote it by L(y^)L($\hat${y}). Our analysis aims at studying the impact of the predictions on the loss, therefore we keep all these other quantities fixed, and vary only the predictions. This is why we only model our loss function $L$ as a function of the predictions $\hat{y}$. In particular, this means that our results typically concern the loss evaluated for a single given data point and a single ground truth label. Of course, we are also often interested in evaluating the loss not only on individual data points, but on batches or entire data sets. Our results on individual data points can then be extended by averaging them, as we discuss further in Section 2.2.3.
@@ 2.2.1 Where do our predictions $\hat{y}_{1},\ldots,\hat{y}_{K}$ come from?
In all the examples presented in the previous paragraph, the $\hat{y}_{1},\ldots,\hat{y}_{K}$ are i.i.d. random variables. Assumptions of this sort will be required for our proofs. Indeed, with the exception of Lemma 1, we will always require our predictions to be random variables with some (weak) dependence structure. We refer to Section 3.3 for a simple example of what can go wrong when there is too much dependency between the individual models.
We want to emphasise that our framework also encompasses situations where we are given a few fully deterministic predictions (for instance, the forecasts of several different physical models). One way to see them as random predictions and to make them amenable to our theory is to assume that the ordering of the $\hat{y}_{1},\ldots,\hat{y}_{K}$ does not matter, or equivalently that we have no knowledge of how these models fare against each other. This can be formalised by assuming that we do not observe $\hat{y}_{1},\ldots,\hat{y}_{K}$ but $\hat{y}_{\sigma(1)},\ldots,\hat{y}_{\sigma(K)}$, where $\sigma$ is a uniformly sampled permutation of $\{1,\ldots,K\}$. Because $\sigma$ is random, $\hat{y}_{\sigma(1)},\ldots,\hat{y}_{\sigma(K)}$ is also random. While not i.i.d. in general, this randomly reordered ensemble satisfies another weaker condition called exchangeability. A sequence of random variables is exchangeable when any of its reorderings has the same distribution, we will give a formal definition in Eq. (13).